Compton Analysis

From PolWiki
Jump to: navigation, search

Compton Analysis Pass Info

The goal of this page is to document all related issues with the analysis of the Compton data from both runs.

Run I

November 2010 to May 2011

RUN II

November 25, 2011 to May 17, 2012

The runlist can be found here: Run_List_-_Run_2

Description/list of EPICS values in Slow_Tree can be found here: Run_2_EPICS

Pass2

Pass 2 officially started on Monday November 5, 2012. A Pass2b had to be started after the discovery of a bug on the analyzer. Pass2b started on Tuesday January 8, 2013.

Here is the status of where run ranges stand (by week)

Week Date Start Run Range Rootfile Status Macro Status Comment
01 11/25/2011 22607 - 22656 In Silo Not started Initial threading of beam through chicane 11/25
02 11/27/2011 22657 - 22789 In Silo Not started
03 12/04/2011 22790 - 22935 In Silo Not started
04 12/11/2011 22936 - 23026 In Silo Not started
05 12/18/2011 23027 - 23094 In Silo Not started
06 12/25/2011 23095 - 23096 -NA- -NA- Winter break shutdown. Test runs only
07 01/01/2012 23097 - 23121 In Silo Not started
08 01/08/2012 23122 - 23153 In Silo Not started
09 01/15/2012 23154 - 23280 In Silo Not started
10 01/22/2012 23281 - 23443 In Silo Not started
11 01/29/2012 23444 - 23601 In Silo Not started
12 02/05/2012 23602 - 23764 In Silo Not started
13 02/12/2012 23765 - 23889 In Silo Not started
14 02/19/2012 23890 - 24051 In Silo Not started
15 02/26/2012 24052 - 24204 In Silo Not started
16 03/04/2012 24205 - 24384 In Silo Not started
17 03/11/2012 24385 - 24526 In Silo Not started
18 03/18/2012 24527 - 24661 In Silo Not started
19 03/25/2012 24662 - 24829 In Silo Not started
20 04/01/2012 24830 - 24972 In Silo Not started
21 04/08/2012 24973 - 25116 In Silo Not started Partial 879MeV running
22 04/15/2012 25117 - 25221 In Silo Not started 879MeV running
23 04/22/2012 25222 - 25223 In Silo Not started 3GeV running, mostly test runs
24 04/29/2012 25224 - 25316 In Silo Not started
25 05/06/2012 25317 - 25541 In Silo Not started
26 05/13/2012 25542 - 25561 In Silo Not started Last week!

Electron detector

Analysis output

The polarization and corresponding CE for most runs in run2 at an early stage of analysis that was shared by Vladas on 24 April 2013. That file is being uploaded here for quick check of the trends through the run2. [link]

The first results from the laser cycle based analysis, without any filtering what-so-ever for the entire run2 can be found at Media:PolAcCE r22659 r25546.txt


DAQ info

Here we list the (non-changing) information related to each run as was recorded on the flags file. For run parameters like pwtl2 etc (which were added during later half of run2), there is a possibility of discrepency between the actual parameter value and that listed here.

The parameter list for the majority of runs during run2 is at Media:RunlistDAQ.txt